OpenStudy (anonymous):

solve using substitution x+y=7 5x+4y=3

OpenStudy (anonymous):

x = 7 - y 5(7 - y) + 4y = 3 35 - 5y + 4y = 3 -y = -32 y = 32 x = 7 - y x = -25 x = -25 y = 32

OpenStudy (anonymous):

see there are two methods of solving these two substitutions given: one is column method and other is that chiaroscuro did. Both of them are approx. same x+y = 7 transposing y to RHS or subtracting both sides by y we will get x = 7-y now putting the value of x on the place of the second substitution: 5(7-y) + 4y = 3 35 - 5y + 4y = 3 35 - y =3 -y = 3 - 35 -y = -32 -1(y) = -32 y = -32/-1 y = 32